OpenStudy (anonymous):

In Linear algebra: The Question gives a system of linear equations: x+ 2y+ 3z+ t = 0 2x+ 4y+ 7z+ 4t = 0 3x+ 6y+ 10z+ 5t = 0 How do you find out the dimension and basis of the solution space? Thanks in advance!

OpenStudy (hba):

Welcome to Open Study.

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

any idea on how to solve?

OpenStudy (hba):

Are you in Uni ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Hi there, you could use gaussian elimination

OpenStudy (anonymous):

Find out how many pivot rows you have, that is equal to it's dimension

OpenStudy (anonymous):

And you can use each independent outcome row as a vector element of an arbitrary basis for the solution set

OpenStudy (anonymous):

OK thanks a lot, I will try and do it now

OpenStudy (anonymous):

when I do the gaussian elimination, I get: 1 2 3 1 0 0 0 -1 -2 0 0 0 0 However, this isn't correct is it?

OpenStudy (anonymous):

Something like that, you have 2 pivot rows there

OpenStudy (anonymous):

1 is a pivot row, what is the other? thanks for all this help btw 0 0 0

OpenStudy (anonymous):

1 2 3 1 is a pivot row 0 0 1 2 is a pivot row too

OpenStudy (anonymous):

To be a pivot, all the elements in the row before the pivot should be 0, the pivot itself should be 1, and everything above the pivot should be 0. But you can always get that 0 by substracting the 2nd pivot row from the first

OpenStudy (anonymous):

OK thanks now that we've got the pivot rows, you then said I can use each independent outcome row as a vector element of an arbitrary basis for the solution set to find the basis I'm a bit confused on what this means?

OpenStudy (anonymous):

Let me see, I don't know the english terms well

OpenStudy (anonymous):

Ok thanks :)- just basically want to know what comes next to find the basis

OpenStudy (anonymous):

Hmm I'm starting to think that I am wrong but we are certainly heading the right way let me recheck something

OpenStudy (anonymous):

hmm well you can solve for x and z now right?

OpenStudy (anonymous):

can you?

OpenStudy (anonymous):

What you can do now is solver for x, y, z, t. You will have 2 'free' variables meaning that the dimension of the solution set is R^4 - 2 = 2. You can thus treat those free variables as scalars and the vectors as a basis

OpenStudy (anonymous):

dim(R^4) - 2*

OpenStudy (anonymous):

Though I can not rigorously prove it, I really need to recheck my LA, but I'm sure the answer would be right

OpenStudy (anonymous):

the answer that my lecturer got is: (x, y, z, t) - y(-2,1,0,0) + t(-2,1,0,0) and (5,0,-2,1) but I am unsure as to how he got this as the basis.

OpenStudy (anonymous):

*** sorry the answer is (x, y, z, t) = y(-2,1,0,0) +t(5,0,-2,1) and the basis is therefore (-2,1,0,0) and (5,0-2,1)

OpenStudy (anonymous):

however, I am not sure how he got that..

OpenStudy (anonymous):

I have the same answer

OpenStudy (anonymous):

So the method is right :)

OpenStudy (anonymous):

Just solve like I told you and if you have questions I'll be here but I need to brb 5 mins

OpenStudy (anonymous):

OK thanks!

OpenStudy (anonymous):

OK so the method is to use gaussian elimination and I get: 1 2 3 1 0 0 0 -1 -2 0 0 0 0 0 0 then from this I can see that the pivot row is 2 = dim(2) after this, how do I solve the above matrix? in order to get the answer that my lecturer gave?

OpenStudy (anonymous):

\[ M*(x y z t)^T = (0 0 0 0)^T \]

OpenStudy (anonymous):

so now you can solve for x y z t

OpenStudy (anonymous):

but you only have two pivot rows which means only two variables can be found, while the other are 'free variables', eg x_i \in R

OpenStudy (anonymous):

how does M∗(xyzt)T=(0000)T help you to solve for xyzt?

OpenStudy (anonymous):

multiply by M-1 or just use the systems of equations approach

OpenStudy (anonymous):

M^-1 *

OpenStudy (anonymous):

I solved it by just using systems of equations

OpenStudy (anonymous):

so when you solve by systems of equations you get? x+2y+3z+t = 0 and -z-2t=0

OpenStudy (anonymous):

and then x=-2t

OpenStudy (anonymous):

sorry z=-2t*

OpenStudy (anonymous):

yep then what is x?

OpenStudy (anonymous):

not sure- how would i get x and y?

OpenStudy (anonymous):

x+2y+3z+t = 0 x = 2y -3(-2t) + t

OpenStudy (anonymous):

y and t are free variables, meaning, they can be any element in \mathbb{R}

OpenStudy (anonymous):

So now show me your solution set for x y z t

OpenStudy (anonymous):

so id they are free variables then we can put t as say 0 therefore, x=2y?

OpenStudy (anonymous):

you could, but then you would get a specific solution, not the solution set

OpenStudy (anonymous):

What you need here is the solution set

OpenStudy (anonymous):

so x = -2y + 5t y = free variable in R (or whatever field you're working in) z = -2t t = free variable in R (or whatever field you're working in)

OpenStudy (anonymous):

So the solution vector space = y*v1 + t*v2

OpenStudy (anonymous):

Try to figure out what vectors v1 and v2 represent

OpenStudy (anonymous):

hmm, they represent the basis but I'm not sure how you would put the solutions into it?

OpenStudy (anonymous):

\[ y * v_1^T + t * v_2^T = ( -2y + 5t , y, -2t, t )^T\]

OpenStudy (anonymous):

that's nearly the answer you should be able to work that out yourself on paper

OpenStudy (anonymous):

Ok so the basis is basically 4 different equations from the matrix

OpenStudy (anonymous):

all put into one solution using the formula of y*v1 + t*v2

OpenStudy (anonymous):

Well yes that's how we did it. This is one way that works but there might be more ways, as is often the case in uni you just have to "work something out yourself" using what you know

OpenStudy (anonymous):

(-2y +5t,y,-2t,t) where did you get the second bit (,y,) from?

OpenStudy (anonymous):

(x y z w) = (...)

OpenStudy (anonymous):

y because y = y ^_^

OpenStudy (anonymous):

oh yes of course haha

OpenStudy (anonymous):

Thank you so much for your help- really appreciated!

OpenStudy (anonymous):

yw